# Visual-RAG: Benchmarking Text-to-Image Retrieval Augmented Generation for Visual Knowledge Intensive Queries

### v1

@@ Appendix E Detailed Results
This section contains additional experimental results. We also include ROUGE-1 recall scores [^9] Implemented with https://pypi.org/project/rouge/, version 1.0.1 for readers who are interested. However, because ROUGE relies on strict word matching, the results for open-ended QA can vary significantly due to the different generation styles across models.
Model Qwen2VL Llama3.2-V GPT-4o
Same Irrel Same Irrel Same Irrel
k=3 40.72 36.27 37.00 36.75 43.90 39.79
5 39.68 37.53 35.92 32.16 45.90 45.03
7 39.36 35.66 36.53 29.55 45.31 40.36
10 39.16 33.30 34.51 27.80 46.14 45.29
15 39.40 35.56 34.59 28.19 46.54 45.64
20 37.73 32.12 - - 45.62 44.50
Table 9: Impact from using highly irrelevant distractor in 1-in-k setting. “Same” denotes hard-negatives from same species, “Irrel” denotes irrelevant images from other species.
Setting Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
Zeroshot 27.57 32.04 26.17 30.21 35.80 47.21 42.78 42.43
Oracle 33.67 44.82 38.38 35.96 43.89 38.45 19.6 33.67
34.26 43.63 37.05 37.25 42.82 39.24 18.49 32.47
30.88 43.82 39.37 39.44 37.12 37.25 18.22 27.69
32.67 40.64 39.97 39.04 45.08 36.25 19.55 33.47
33.07 43.03 38.11 37.25 38.65 37.05 16.94 33.67
AVG 32.91 43.19 38.58 37.79 41.51 37.65 18.56 32.19
STD 1.285 1.563 1.136 1.433 3.449 1.188 1.096 2.566
Non-clue 26.69 33.27 31.21 33.27 30.28 14.74 4.46 8.17
27.49 34.26 34.06 35.86 34.66 17.73 7.41 8.76
30.28 33.86 32.07 37.05 29.88 17.93 6.71 10.16
27.09 33.47 29.68 34.66 32.53 16.53 7.07 8.96
27.29 31.47 30.61 33.86 30.81 16.14 5.59 8.57
AVG 27.77 33.27 31.53 34.94 31.63 16.61 6.25 8.92
STD 1.435 1.073 1.663 1.527 1.972 1.296 1.211 0.750
@@
1 in K Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
k=3 32.27 40.44 37.58 37.45 39.04 41.63 42.94 32.27
32.27 39.64 36.52 35.26 34.79 44.22 49.09 32.67
34.06 40.24 37.71 37.85 34.53 44.02 47.43 35.26
33.86 41.63 36.12 39.64 38.84 43.03 41.99 30.68
32.67 41.63 35.65 35.86 37.78 46.61 46.75 33.47
AVG 33.03 40.72 36.72 37.21 37.00 43.90 45.64 32.87
STD 0.871 0.885 0.903 1.732 2.187 1.828 3.039 1.678
k=5 33.86 39.84 34.39 40.93 37.78 46.81 47.8 37.25
33.20 39.64 34.46 38.05 33.20 46.02 46.54 34.86
34.46 40.24 35.45 44.2 35.06 47.01 46.61 36.06
33.67 39.64 35.99 36.85 37.12 44.42 43.61 35.26
34.66 39.04 35.79 38.45 36.45 45.22 47.18 36.45
AVG 33.97 39.68 35.22 39.70 35.92 45.90 46.35 35.98
STD 0.594 0.434 0.748 2.923 1.825 1.087 1.613 0.950
k=7 31.08 38.65 35.45 34.66 37.91 43.03 46.25 37.4
32.27 39.44 34.13 37.05 35.39 44.02 44.97 35.88
32.27 38.84 36.32 33.47 34.59 45.29 44.71 36.83
36.85 39.44 35.06 36.65 35.85 46.22 45.66 35.88
34.26 40.44 33.06 36.46 38.91 48.01 45.71 37.02
AVG 33.35 39.36 34.80 35.66 36.53 45.31 45.46 36.60
STD 2.267 0.699 1.253 1.530 1.809 1.935 0.618 0.690
k=10 32.27 39.04 35.72 37.00 33.53 45.62 46.73 34.15
33.07 39.64 37.05 37.10 34.45 44.82 48.79 37.65
32.87 39.84 35.12 37.70 36.39 46.41 49.66 38.25
31.47 38.45 35.78 36.23 33.20 47.61 43.54 39.24
31.67 38.84 34.39 34.82 34.99 46.22 45.76 38.84
AVG 32.27 39.16 35.61 36.57 34.51 46.14 46.90 37.63
STD 0.707 0.573 0.980 1.109 1.269 1.031 2.440 2.034
k=15 31.27 40.44 33.86 42.01 31.94 45.42 47.6 34.15
33.86 39.24 35.99 36.55 34.92 48.01 49.63 37.79
30.08 40.24 34.66 36.38 35.52 47.81 42.42 38.74
32.47 38.84 36.65 36.38 36.58 46.02 47.6 39.5
33.47 38.25 33.79 35.27 34.00 45.42 45.66 39.12
AVG 32.23 39.40 34.99 37.32 34.59 46.54 46.58 37.86
STD 1.565 0.929 1.283 2.672 1.754 1.280 2.717 2.169
k=20 33.47 38.05 33.93 - - 44.82 51.15 35.67
29.48 39.44 35.32 - - 44.62 47.99 34.76
33.47 36.25 31.94 - - 45.22 47.13 36.39
32.47 37.65 34.00 - - 47.21 48.57 33.47
33.07 37.25 35.19 - - 46.22 53.01 36.85
AVG 32.39 37.73 34.08 - - 45.62 49.57 35.43
STD 1.679 1.168 1.358 - - 1.083 2.438 1.350
@@
Setting Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
Zeroshot 14.66 14.87 17.03 20.44 14.05 11.96 12.58 9.65
Oracle 15.26 26.75 17.16 25.83 6.72 16.48 16.35 24.59
15.85 26.63 19.01 25.96 8.1 15.9 14.1 23.26
15.22 25.86 18.72 29.25 7.19 18.44 14.77 23.63
15.79 25.08 19.46 28.42 10.39 20.07 14.58 23.83
15.27 26.44 17.06 25.86 7.88 20.36 11.47 26.57
AVG 15.48 26.15 18.28 27.06 8.06 18.25 14.25 24.38
STD 0.313 0.690 1.102 1.644 1.416 2.028 1.771 1.319
Non-clue 9.49 21.59 8.07 23.25 3.95 9.23 9.04 17.11
9.27 21.13 9.24 23.94 5.4 7.86 10.03 17.18
8.89 20.96 10.78 24.4 2.87 10.06 9.21 17.42
9.85 21.23 8.69 24.63 5 6.62 11.4 15.26
9.27 20.86 9.05 24.18 4.05 9.9 10.14 16.09
AVG 9.35 21.15 9.17 24.08 4.25 8.73 9.96 16.61
STD 0.351 0.283 1.006 0.530 0.990 1.466 0.938 0.912
Table 12: Baselines rouge scores for reference. Please note that rouge score fluctuates largely across models, as the generated text style largely affect the matching with ground truth.
1 in K Phi3.5-V Qwen2VL InternVL2.5 Pixtral Llama3.2-V GPT-4o Gemini Claude
k=3 10.49 23.80 8.79 24.31 6.92 26.23 16.14 23.27
10.30 25.47 9.65 24.12 6.64 20.81 18.71 22.65
10.57 23.89 8.22 25.83 5.55 24.87 19.38 22.69
9.89 24.72 8.78 25.27 7.97 23.78 17.91 22.74
10.29 25.80 9.34 25.79 6.76 24.18 19.19 21.31
AVG 10.31 24.74 8.96 25.06 6.77 23.97 18.27 22.53
STD 0.263 0.903 0.554 0.809 0.861 1.999 1.317 0.728
k=5 12.18 25.82 7.77 25.82 8.71 22.44 16.94 22.37
9.87 23.39 8.22 25.45 4.79 21.04 16.36 23.2
10.50 23.54 8.65 26.76 5.55 25.28 16.95 23.62
11.76 22.36 7.72 22.11 7.65 17.79 16.17 23.17
10.30 24.67 8.62 24.56 5.12 28.04 17.84 24.35
AVG 10.92 23.72 8.20 24.94 6.36 22.92 16.85 23.34
STD 0.995 0.967 0.446 1.768 1.721 3.933 0.652 0.722
k=7 9.28 23.26 8.63 21.65 7.57 22.62 15.67 24.38
10.72 24.28 8.17 24.61 5.66 25.55 20.08 21.73
11.45 24.74 8.38 21.51 7.14 20.67 16.35 24
12.10 24.81 8.22 24.40 7.63 25.97 17.62 22.79
12.48 24.32 7.67 23.57 7.01 24.4 18.82 23.07
AVG 11.21 24.28 8.21 23.15 7.00 23.84 17.71 23.19
STD 1.268 0.619 0.353 1.484 0.796 2.197 1.794 1.046
k=10 11.85 23.90 7.65 25.71 5.81 23.86 18.92 23.03
12.15 24.00 8.81 25.50 6.93 24.44 17.5 21.05
11.83 23.69 8.19 24.96 6.15 23.38 18.64 21.52
11.79 24.45 8.77 25.28 6.62 24.34 17.57 23.31
11.06 24.63 8.06 26.36 6.78 26.39 19.17 21.71
AVG 11.74 24.13 8.30 25.56 6.46 24.48 18.36 22.12
STD 0.404 0.392 0.493 0.525 0.466 1.147 0.776 0.990
k=15 11.42 23.21 8.31 24.68 7.44 21.5 19.12 20.55
11.52 24.76 8.93 25.77 8.41 25.05 19.79 21.93
10.14 24.49 7.84 26.46 7.07 23 18.47 24.1
11.77 23.49 8.11 23.52 7.89 28.41 20.25 24.11
11.77 23.34 8.52 26.43 6.96 21.25 20.24 23
AVG 11.32 23.86 8.34 25.37 7.55 23.84 19.57 22.74
STD 0.680 0.714 0.414 1.262 0.601 2.968 0.770 1.520
k=20 13.09 22.95 8.02 - - 24.21 18.57 20.79
11.51 23.58 9.40 - - 21.31 16.66 20.56
11.76 23.24 8.30 - - 22.27 17.95 23.95
11.85 23.94 8.20 - - 21.99 19.25 20.61
11.95 22.80 9.00 - - 25.61 20.19 23.31
AVG 12.03 23.30 8.58 - - 23.08 18.52 21.84
STD 0.614 0.465 0.588 - - 1.779 1.333 1.648

### v2

@@ Appendix E Detailed Results
This section contains detailed experimental results. The breakdown results for main experiments can be found in Table 13, 15. We also include ROUGE-1 Recall scores [^8] Implemented with https://pypi.org/project/rouge/, version 1.0.1 for readers who are interested in Table 14 and 16. However, ROUGE relies on strict word matching, the results for open-ended QA can vary significantly due to the different generation styles across models.
Setting Phi4-$M$ Qwen2.5VL InternVL3 Pixtral Llama3.2-V GPT-4o Gemini Claude
Zeroshot 35.16 38.90 39.17 41.71 32.35 53.74 60.43 54.28
Non-clue 34.76 30.08 29.28 42.74 40.37 14.97 17.11 21.39
GT clue 45.05 42.51 42.11 46.66 46.26 62.17 62.17 55.35
44.92 38.50 41.44 47.86 45.99 56.95 61.76 60.56
45.94 43.58 43.32 47.73 47.46 60.16 63.10 55.88
46.79 43.72 44.92 45.19 50.40 60.29 64.57 55.88
42.51 40.64 46.66 48.13 48.93 59.49 62.79 56.28
AVG 45.04 41.79 43.69 47.11 47.81 59.81 62.88 56.79
STD 1.604 2.213 2.123 1.212 1.857 1.884 1.081 2.133
@@
Setting Phi4-$M$ Qwen2.5VL InternVL3 Pixtral Llama3.2-V GPT-4o Gemini Claude
Zeroshot 21.46 27.06 25.66 25.88 16.63 24.00 26.47 35.38
Non-clue 13.96 12.33 19.81 23.52 3.96 13.76 16.01 24.14
GT clue 18.79 27.78 20.64 28.11 15.55 25.25 24.34 32.36
19.29 25.89 21.16 26.74 15.73 26.51 22.44 33.88
18.74 27.43 21.36 27.68 16.18 25.64 22.94 33.45
19.70 27.49 22.71 26.64 16.27 24.26 24.15 32.37
18.91 26.89 24.19 26.48 14.53 26.67 24.38 33.46
AVG 19.09 27.10 22.01 27.13 15.65 25.67 23.65 33.10
STD 0.405 0.747 1.437 0.721 0.696 0.984 0.898 0.697
Top-K Retrieval-Augmented Generation
k=1 16.03 13.58 21.71 23.58 4.54 14.48 20.57 28.22
k=3 16.92 7.95 20.66 22.32 7.37 19.60 24.50 32.78
k=5 14.99 11.45 18.98 21.92 5.46 22.59 25.67 34.04
k=7 16.15 12.81 17.62 21.51 6.11 24.19 27.92 34.73
k=10 16.61 13.88 18.40 22.49 6.73 23.49 28.17 35.18
k=15 16.55 13.55 16.94 24.60 5.72 23.74 31.12 35.43
k=20 16.75 13.51 14.93 - - 25.50 30.30 37.95
Table 14: Baselines and Top-K RAG ROUGE scores for reference. Please note that ROUGE score fluctuates largely across models, as the generated text style largely affect the matching with ground truth.
1 in K Phi4-$M$ Qwen2.5VL InternVL3 Pixtral Llama3.2-V GPT-4o Gemini Claude
k=3 41.04 46.93 39.44 48.26 46.26 47.99 59.09 48.26
43.54 47.46 41.67 46.12 46.75 50.22 60.29 48.80
40.78 45.05 41.18 48.80 48.53 47.33 58.82 47.46
41.98 46.26 43.85 45.86 47.73 50.27 60.29 47.33
42.25 48.53 39.17 46.52 44.39 48.93 59.36 48.40
AVG 41.92 46.85 41.06 47.11 46.73 48.95 59.57 48.05
STD 1.097 1.303 1.895 1.329 1.576 1.314 0.684 0.632
k=5 43.45 48.13 37.97 46.39 47.59 54.41 61.50 50.00
39.30 48.53 40.37 47.59 45.05 51.28 59.49 49.06
38.50 47.55 36.32 46.52 43.72 54.28 62.57 49.06
39.44 49.06 38.24 46.66 45.86 52.67 59.36 50.40
39.97 46.93 39.30 46.26 43.72 53.48 60.56 50.80
AVG 40.13 48.04 38.44 46.68 45.19 53.22 60.70 49.86
STD 1.928 0.831 1.518 0.528 1.624 1.292 1.362 0.787
k=7 40.37 48.80 40.64 46.39 43.32 56.68 61.93 51.20
41.04 47.06 36.36 47.59 42.51 56.37 61.90 52.81
39.57 47.46 40.64 46.52 46.79 54.41 61.36 50.48
40.64 48.26 41.44 46.66 42.38 51.34 61.76 50.94
42.38 47.33 39.71 46.26 45.99 58.02 60.70 49.60
AVG 40.80 47.78 39.76 46.68 44.20 55.36 61.53 51.01
STD 1.034 0.724 1.996 0.528 2.053 2.593 0.517 1.178
k=10 40.24 45.19 37.43 44.65 41.44 53.61 62.70 50.27
41.00 47.06 39.44 46.12 37.17 57.49 60.56 51.07
37.97 48.66 38.64 44.79 42.38 56.23 63.90 53.88
39.57 49.06 38.50 44.61 42.91 55.75 63.64 53.21
36.63 47.19 38.50 43.98 45.59 58.16 62.70 50.00
AVG 40.80 47.43 38.50 44.83 41.90 56.25 62.70 51.69
STD 1.034 1.531 0.716 0.786 3.060 1.761 1.314 1.758
k=15 41.31 51.34 39.30 44.92 43.32 57.62 64.17 52.01
39.57 49.73 37.83 43.72 39.17 55.88 64.71 51.20
39.71 51.07 38.90 43.58 38.50 55.08 63.90 54.68
42.25 47.59 39.17 43.54 42.11 57.49 62.37 54.81
39.17 48.13 40.51 43.98 42.51 55.75 64.88 50.67
AVG 39.08 49.57 39.14 43.95 41.12 56.36 64.01 52.67
STD 1.768 1.688 0.959 0.570 2.146 1.130 0.997 1.950
k=20 41.58 47.33 39.30 - - 56.95 65.37 52.54
39.84 49.33 37.03 - - 57.22 64.57 50.13
41.18 50.94 39.44 - - 54.81 62.03 53.21
39.71 47.15 42.78 - - 55.48 62.43 51.60
41.84 47.99 36.90 - - 58.02 62.97 53.21
AVG 37.25 48.55 39.09 - - 56.50 63.47 52.14
STD 0.992 1.588 2.389 - - 1.316 1.434 1.302
@@
1 in K Phi4-$M$ Qwen2.5VL InternVL3 Pixtral Llama3.2-V GPT-4o Gemini Claude
k=3 16.93 7.05 19.39 24.89 6.12 20.61 28.82 35.80
16.88 7.43 21.30 25.07 6.23 22.40 29.24 35.15
16.55 6.37 20.01 25.09 6.70 20.58 27.82 32.25
17.34 7.24 22.95 23.82 5.70 20.96 28.97 33.58
15.25 7.54 19.71 24.59 5.66 21.69 29.18 35.05
AVG 16.59 7.13 20.67 24.69 6.08 21.25 28.81 34.37
STD 0.800 0.462 1.466 0.527 0.427 0.784 0.576 1.435
k=5 17.55 9.72 19.33 24.35 5.63 24.47 27.78 35.70
15.90 10.50 18.26 24.85 7.20 22.47 29.19 36.12
15.54 11.02 15.34 24.03 6.12 21.93 30.06 33.99
16.46 11.55 20.07 23.60 5.90 22.01 28.95 34.39
16.22 10.36 19.00 23.37 5.86 23.01 28.13 34.48
AVG 16.33 10.63 18.40 24.04 6.14 22.78 28.82 34.94
STD 0.763 0.692 1.830 0.591 0.616 1.039 0.901 0.920
k=7 14.37 12.70 17.83 23.93 7.25 24.52 29.94 34.47
14.28 12.09 16.69 24.10 6.11 23.09 28.13 34.93
15.18 10.95 18.25 23.60 6.09 24.05 29.74 32.26
15.65 11.51 17.68 23.81 6.04 23.47 30.29 35.47
16.17 11.99 18.22 22.65 5.78 23.77 31.00 35.81
AVG 15.13 11.85 17.73 23.62 6.25 23.78 29.82 34.59
STD 0.815 0.657 0.633 0.571 0.572 0.546 1.060 1.398
k=10 16.04 12.20 15.97 22.10 6.06 21.83 31.15 34.62
14.90 12.62 15.92 22.78 6.11 22.11 30.38 35.07
15.24 13.65 16.39 23.42 5.64 24.25 30.47 35.48
15.55 12.46 16.59 23.37 5.39 22.95 32.16 34.23
15.56 12.26 17.49 23.85 7.08 25.03 29.40 33.65
AVG 15.46 12.64 16.47 23.10 6.06 23.23 30.71 34.61
STD 0.423 0.590 0.635 0.678 0.646 1.376 1.022 0.713
k=15 15.43 12.39 17.70 26.02 5.51 25.42 31.44 34.84
15.26 13.18 16.60 25.30 7.23 24.87 31.85 35.15
15.22 12.76 14.40 25.35 5.49 24.45 30.32 34.60
15.32 12.57 15.20 25.10 5.72 24.37 31.54 35.83
14.91 12.79 16.12 26.04 5.76 24.90 32.22 35.69
AVG 15.23 12.74 16.00 25.56 5.94 24.80 31.47 35.22
STD 0.195 0.295 1.271 0.437 0.730 0.420 0.713 0.531
k=20 16.24 12.60 15.07 - - 24.29 31.76 36.16
15.03 13.46 14.38 - - 23.72 31.80 34.58
16.00 13.64 15.57 - - 23.43 30.94 36.53
15.85 13.42 15.13 - - 24.37 31.47 34.87
15.40 13.03 14.35 - - 24.09 31.84 35.40
AVG 15.70 13.23 14.90 - - 23.98 31.56 35.51
STD 0.486 0.417 0.525 - - 0.397 0.377 0.829
@@
Appendix F Qualitative Error Case Analysis
Providing the ground-truth clue images effectively makes the problem a standard VQA task, which is considered to be “solved” for the state-of-the-art proprietary models. However, our experiment results shown that their improvements over zero-shot baseline is marginal. We present a few typical error cases from proprietary models under GT clue setting in Table 17.
Question: When resting, Acadian Hairstreak (scientific name: Satyrium acadica) usually fold its wings and the abodmen is hidden; what color is its abdomen?
Ground Truth Answer: [white, light grey]
Gemini’s Answer: In the provided image, the abdomen of the Acadian Hairstreak is obscured by its folded wings, making its color not visible.
Question: What color is the stamen tip inside the bell-shaped flower of western moss-heather (scientific name: Cassiope mertensiana)? (Explanation: opening of the flower usually points downward)
Ground Truth Answer: light yellow
Gemini’s Answer: Red
Table 17: Sample cases that Gemini failed to answer even with ground-truth clue image. Fine-grained details in the image of area of few pixels remain challenging even for the state-of-the art proprietary models.